OpenStudy (beacandy):

F a n + M e d a l The length of a rectangle is 3 units shorter than one-third of the width, x. Enter an expression in the box that represents the perimeter of the rectangle. ______

OpenStudy (shadow_helper):

first write out how you would... and ill see if its right or wrong.

OpenStudy (mazmaz1):

L =(x/3)-3

OpenStudy (shadow_helper):

y=-3+3/x

OpenStudy (mazmaz1):

lenth is 3less so -3 than one thir of width so x/3

OpenStudy (saadsword):



OpenStudy (shadow_helper):

that^^^... i didnt wanted to write the whole thing down too...

OpenStudy (mazmaz1):

me and sword got same answer just written in different format mine is using l instead of y ans I have the order right

OpenStudy (saadsword):

Perimeter=2(3-1/3 (w)) +2(w)

OpenStudy (saadsword):

simplify

OpenStudy (beacandy):

Thanks guys

OpenStudy (mazmaz1):

your welcome